# From Age Estimation to Age-Invariant Face Recognition: Generalized Age Feature Extraction Using Order-Enhanced Contrastive Learning

### v1

@@ IV-B3 Implementation Details
TABLE IV: Comparison of cross-dataset results, where MO represents the MORPH II dataset, FG represents the FG-NET dataset, and AG represents the AgeDB dataset. Cls indicates conventional classification-based methods, and Con indicates contrastive learning-based methods. The results from methods used for comparison in this table are produced using officially released code. The best results are highlighted in bold, and the second-best results are underlined.
@@
IV-C Age Estimation Results
IV-C1 Homogeneous-Dataset Results
The homogeneous-dataset results for the MORPH II and FG-NET datasets are presented in Tables II and III, respectively. In Table II, results for all three settings are reported. As shown, OrdCon achieves either state-of-the-art or comparable results across all settings. Despite not using the IMDB-WIKI dataset for pretraining, our method outperforms those that do, highlighting the importance of modeling age progression between samples, whereas most methods only learn a direct mapping between features and labels. Notably, our method also outperforms RNC [51], which uses contrastive learning to rank the order among samples. By incorporating metric learning with soft proxy matching constraints, our method learns features that are positioned closer to the center of each age, resulting in a more compact feature distribution aligned with natural age progression. Fig. 4 illustrates how the learned age feature space on the MORPH II dataset evolves during training. It also highlights the dataset’s imbalance, with the majority of younger faces represented by the thicker dark green end and the minority of older faces represented by the lighter green end.
In Table III, the homogeneous-dataset results on the FG-NET dataset show a similar trend to those on the MORPH II dataset. However, due to the small size of the FG-NET dataset, methods pretrained with the IMDB-WIKI dataset achieve relatively lower MAE compared to those trained from scratch. Despite this, OrdCon achieves state-of-the-art results among methods without additional pretraining, thanks to the extra samples leveraged by our approach.
@@ IV-C2 Cross-Dataset Results
For cross-dataset experiments, we use all three datasets MORPH II, FG-NET, and AgeDB to form six training-testing pairs. Specifically, we train the model on one dataset and test it on one of the other two. The results are presented in Table IV. As shown, the comparison includes three conventional regression-based methods and four contrastive learning-based methods, including OrdCon. For each dataset, we use the entire dataset for either training or testing.
@@
Since SimCLR and SupCon are designed for classification tasks, their MAEs are significantly higher compared to regression-targeted contrastive learning methods like RNC and OrdCon, especially when the age label is not used in SimCLR. Both RNC and OrdCon achieve lower MAE values compared to conventional regression-based methods, owing to the superior generalization ability of contrastive learning [91].
We compare the learned feature spaces of GOL, SupCon, RNC, and OrdCon in Fig.5. The features are extracted from facial images in the FG-NET dataset using a model trained on the AgeDB dataset. Facial images from two sample identities are also shown at the bottom of the figure. Since SupCon is not designed for regression tasks, its learned feature space lacks a clear indication of age progression. In contrast, the feature spaces learned by GOL and RNC show a clear trend from younger to older faces. However, for some sample identities, the aging process does not always align with the general progression.
GOL uses similar techniques to OrdCon, but without proxy-based metric learning and the soft proxy matching loss, features are pushed equally away from all other age clusters, which can lead to misplacement. While RNC considers age differences among features, it does not take ordinal information into account, resulting in the absence of a clear aging direction. In comparison, the feature space learned by OrdCon shows a more distinct aging progression. Although some features are misplaced, the aging process of the sample identities is well aligned with the general progression learned by OrdCon.
Figure 5: Comparison of learned age feature spaces from GOL, SupCon, RNC, and OrdCon, along with images of two sample identities. Best viewed in color.
@@ IV-D1 Homogeneous-Dataset Results
Similar to the homogeneous-dataset experiments on age estimation, we conduct experiments on the MORPH II and FG-NET datasets. In addition to our preliminary work [25], CACon, we also include two commonly used contrastive learning methods, SimCLR and SupCon, for comparison to highlight the differences between our method and other contrastive approaches. The results on the two datasets are presented in Tables V and VI. As shown in Table V, OrdCon achieves the best results in both settings among contrastive learning-based methods and comparable results to state-of-the-art AIFR methods. Table VI shows that our method outperforms other AIFR methods on the FG-NET dataset across all three settings. Since OrdCon synthesizes one sample per age group for each facial image during training, it is less affected by the limited number of training samples in the FG-NET dataset. It also marginally outperforms CACon, which synthesizes only one additional sample at a random age group.
@@ IV-D2 Cross-Dataset Results
Since there is limited work reporting results on cross-dataset AIFR experiments, for fair comparison, we only include methods with officially released code, including three conventional classification-based methods and three contrastive learning-based methods. Again, in cross-dataset experiments, we use the entire dataset for either training or evaluation. The results are presented in Table VII. Due to the disentangling of identity features from robust and generalized age features in our method, OrdCon outperforms all conventional classification-based methods as well as contrastive learning-based methods that do not model natural age progression. Among conventional classification-based methods, IEFP achieves results most comparable to ours. Notably, IEFP and OrdCon share a similar strategy—learning age features first and then disentangling them from identity features.
@@
TABLE VIII: Evaluation of the contribution of each component in OrdCon on the AgeDB dataset for both age estimation and AIFR tasks.
Method MAE Rank-1
contrast (SupCon) 9.32 91.03
contrast + Order 6.74 94.87
contrast + Metric 8.55 95.04
contrast + Order + Metric 6.15 97.21
contrast + Order + Soft Metric (OrdCon) 5.95 97.44
IV-E Ablation Study
@@
Figure 6: Comparison of age feature spaces on the AgeDB test set using different combinations of order learning and metric learning. Metric learning refers to proxy-based metric learning with soft proxy matching loss.
IV-E1 Contribution of Each Component
@@
We select SupCon [22] as the baseline, as it represents the fundamental form of supervised contrastive learning. In addition to the baseline model, we design three variations. The first two variations use either the order learning objective or the metric learning objective independently to demonstrate their effectiveness compared to SupCon with a conventional contrastive learning objective. Next, we combine order learning with hard metric learning using Eq. (13) to make the distribution of each age more compact, thereby improving performance. Finally, we replace hard metric learning with its soft counterpart to create the complete OrdCon, making the metric learning constraint proportional to age differences among samples for better modeling of age progression.
@@
Fig. 6 compares the age feature spaces learned on the AgeDB test set using different combinations of order learning and metric learning. The age progression becomes increasingly fine-grained as more objectives are added to the model. Notably, metric learning with the soft proxy matching loss alone can also capture age progression to some extent, although the features representing different ages are not as well separated compared to the feature spaces learned with order learning.
When using SupCon for AIFR, and to maintain consistency with the gradient reversal strategy in OrdCon, we replace the order learning objective and metric learning loss with an L1 loss. The results, presented in Table VIII, show significant improvements in performance on both tasks when order learning and metric learning are added to the contrastive learning model. The performance is further refined when relative distance differences are incorporated into the metric learning objective.
@@
TABLE IX: Comparison of Rank-1 accuracy under different age group granularities on the AgeDB dataset.
@@ IV-E2 Age Group Granularity
For AIFR, we also test how different age group partitions affect accuracy. Since facial aging effects are subtle over a few years and to control the computational complexity of our method, we test age group granularities ranging from 5 to 10 years. The results for different age group granularities on the AgeDB dataset are presented in Table IX. The best results are achieved when the entire dataset is divided into groups with an age span of 6 years (e.g., 0-5, 6-11, etc.). An age span of 5 years also delivers similar results but with higher computational complexity, as more samples need to be synthesized and used for contrastive learning. As the age span in each group increases, the AIFR accuracy gradually decreases due to higher intra-class dynamics.
V Conclusion
@@
In this paper, we proposed OrdCon to extract generalized age features for both age estimation and AIFR tasks. Our method leverages contrastive learning with order learning to model natural age progression, along with metric learning to minimize intra-class variance within each age cluster. The proposed method can be directly applied to age estimation. For AIFR, we adopt a multitask feature extractor to simultaneously extract age and identity features, with a GRL used to make age information indistinguishable, achieving age-invariant identity features. By utilizing contrastive learning for better generalization capability and order learning along with metric learning for robust age feature modeling, our model achieves either state-of-the-art or comparable performance in homogeneous-dataset experiments for both age estimation and AIFR task and outperforms other methods in cross-dataset experiments.
VI Acknowledgment

### v2

@@ IV-C1 Homogeneous-Dataset Results
TABLE IV: Comparison of cross-dataset results, where MO represents the MORPH II dataset, FG represents the FG-NET dataset, and AG represents the AgeDB dataset. Reg indicates conventional regression-based methods, and Con indicates contrastive learning-based methods. The results from the methods used for comparison in this table are produced using officially released code. The best results are highlighted in bold, and the second-best results are underlined.
@@ IV-C2 Cross-Dataset Results
For cross-dataset experiments, we utilize all three datasets, MORPH II, FG-NET, and AgeDB, to form six training-testing pairs. Specifically, we train the model on one dataset and test it on one of the other two. The results are presented in Table IV. As shown, the comparison includes three conventional regression-based methods and seven contrastive learning-based methods, including OrdCon. For each dataset, we use the entire dataset for either training or testing.
@@
Since SimCLR and SupCon are designed for classification tasks, their MAEs are significantly higher compared to regression-targeted contrastive learning methods, such as RNC and OrdCon. Both achieve lower MAE values compared to conventional regression-based methods, owing to the superior generalizability of contrastive learning [19]. It is worth noting that OrdCon outperforms MCGRL in all the scenarios, highlighting the importance of modeling the aging progression for improving the model’s generalizability.
We compare the learned feature spaces of GOL, SupCon, RNC, and OrdCon in Fig.5. The features are extracted from facial images in the FG-NET dataset using a model trained on the AgeDB dataset. Facial images from two sample identities are also shown at the bottom of the figure. Since SupCon is not designed for regression tasks, its learned feature space lacks a clear indication of age progression. In contrast, the feature spaces learned by GOL and RNC show a clear trend from younger to older faces. However, for some sample identities, the aging process does not always align with the general progression of aging.
GOL uses similar techniques to OrdCon, but without proxy-based representation learning and the soft proxy matching loss, features are pushed equally away from all other age clusters, which can lead to misplacement. While RNC considers age differences among features, it does not account for ordinal information, resulting in the absence of a clear aging direction. In comparison, the feature space learned by OrdCon shows a more distinct aging progression.
Figure 5: Comparison of learned age feature spaces from four methods, along with images of two sample identities. Best viewed in color and zoomed in.
@@ IV-D1 Homogeneous-Dataset Results
Similar to the homogeneous-dataset experiments on age estimation, we conduct experiments on the MORPH II and FG-NET datasets. In addition to our preliminary work [26], CACon, we also include two commonly used contrastive learning methods, SimCLR and SupCon, for comparison to highlight the differences between our method and other contrastive approaches. The results on the two datasets are presented in Tables V and VI. As shown in Table V, OrdCon achieves the best results among contrastive learning-based methods in both settings and comparable results to state-of-the-art AIFR methods. However, it underperforms methods like IEFP [54], which leverages carefully designed purification units to learn dataset-specific correlation between age and identity features. Table VI shows that our method outperforms other AIFR methods on the FG-NET dataset across all three settings. Since OrdCon leverages one sample per age group for each facial image during training, it is less affected by the limited number of training samples in the FG-NET dataset. It also outperforms CACon, which uses only one additional sample at a random age group.
@@ IV-D2 Cross-Dataset Results
Since there is limited work reporting results on cross-dataset AIFR experiments, for fair comparison, we only include methods with officially implementations, including three conventional classification-based methods and three contrastive learning-based methods. Again, in cross-dataset experiments, we use the entire dataset for either training or evaluation. The results are presented in Table VII. Due to the disentangling of identity features from robust and generalized age features in our method, OrdCon outperforms all conventional classification-based methods as well as contrastive learning-based methods that do not model natural age progression. Among conventional classification-based methods, IEFP achieves results most comparable to ours. Notably, IEFP and OrdCon share a similar strategy—learning age features first and then disentangling them from identity features.
IV-E Generalizability Evaluation
Apart from task-specific cross-dataset experiments, we also evaluate the generalizability of our framework against other contrastive learning frameworks using the alignment of positive samples and the divergence of direction vectors. For both matrices, we use the MOPRH II dataset as the training set, as it is the largest among the three we are using, and use the FG-NET dataset as the test set. As shown in Table VIII, OrdCon achieves the best results on both matrices. In terms of the alignment of positive samples, OrdCon outperforms CACon due to its aging progression modeling and the use of multiple augmented samples at different age groups during training. Furthermore, both achieve smaller alignment compared to general frameworks like SimCLR and SupCon. Regarding the divergence of direction vectors, OrdCon outperforms RNC because the latter only considers absolute age differences among samples and overlooks the ordinal relationships.
TABLE VIII: Evaluation of the generalizability for state-of-the-art methods and OrdCon.
[…]
SimCLR [23] 0.77 -
SupCon [24] 0.63 -
CACon [26] 0.47 -
RNC [25] - 0.41
CLOC [36] - 0.28
MCGRL [80] - 0.32
OrdCon (ours) 0.40 0.56
@@ IV-E Generalizability Evaluation
TABLE IX: Evaluation of the contribution of each component in OrdCon on the AgeDB dataset for both age estimation and AIFR tasks.
[…]
9.32 91.03 0.63 -
6.74 94.87 - 0.42
hard 8.55 95.04 0.46 -
hard 6.15 97.21 0.42 0.53
soft 5.95 97.44 0.40 0.56
@@
Figure 6: Comparison of age feature spaces on the AgeDB test set using different combinations of order learning and proxy-based learning.
IV-F Ablation Study
IV-F1 Contribution of Each Component
@@
We select SupCon [24] as the baseline as it has only a contrastive component and represents the fundamental form of supervised contrastive learning. In addition to this baseline model, we design three variations. The first two variations use either the order learning-based contrastive objective or the proxy-based contrastive objective independently to demonstrate their effectiveness compared to SupCon with a conventional contrastive learning objective. Next, we combine order learning with hard proxy-based learning using Eq. (13) to make the distribution of each age more compact, thereby improving performance. Finally, we replace hard proxy-based learning with its soft counterpart to create the complete OrdCon, making the proxy-based constraint proportional to age differences among samples for better modeling of age progression.
Although all the frameworks in Table IX can be used for age estimation and AIFR to some extent, similar to Table VIII, we only evaluate alignment of positive samples on frameworks targeted on classification tasks, i.e., frameworks without the order learning component. Likewise, we only evaluate divergence of direction vectors on frameworks with the order learning component for regression tasks.
@@
Fig. 6 compares the age feature spaces learned on the AgeDB test set using different combinations of order learning and proxy-based learning. The age progression becomes increasingly fine-grained as more objectives are added to the model. Notably, soft proxy matching loss alone can also roughly capture age progression. However, the features representing different ages are not as well separated as those learned with order learning.
@@
TABLE X: Comparison of Rank-1 accuracy under different age group granularities on the AgeDB dataset.
@@ IV-F2 Age Group Granularity
For AIFR, we also test how different age group partitions affect accuracy. Since facial aging effects are subtle over a few years and to control the computational complexity of our method, we test age group granularities ranging from 5 to 10 years. The results for different age group granularities on the AgeDB dataset are presented in Table X. The best results are achieved when the entire dataset is divided into groups with a 6-year age span (e.g., 0-5, 6-11, etc.). An age span of 5 years also yields similar results, albeit with higher computational complexity, as more samples need to be synthesized and utilized for contrastive learning. As the age span in each group increases, the AIFR accuracy gradually decreases due to higher intra-class dynamics.
V Discussion
V-A Applicability
While the current work focuses on age feature extraction for two mainstream age-related tasks, i.e., age estimation and AIFR, the underlying principles of OrdCon could be extended to other age-related tasks and facial attributes.
V-A1 To other age-related tasks
Beyond age estimation and AIFR, OrdCon holds significant potential for other notable age-related tasks such as age synthesis and age privacy.
For age synthesis, the generalized age features extracted by OrdCon, which are specifically designed to capture the natural progression of aging, could guide fine-grained face aging synthesis. As highlighted by Yao et al. [55], robust age representation is crucial for accurate age synthesis. By learning a feature space where age transitions are smooth and well-represented, OrdCon could guide generative models to produce more accurate facial images with rich aging features. More specifically, our learned age features, structured along an aging trajectory, could provide strong conditional input for GAN-based models or serve as a guide for latent space manipulation in diffusion models.
Regarding age privacy, the concept of extracting and potentially disentangling age features also has significant implications. Zhang et al. [89] discuss methods for concealing attributes while retaining utility. Similarly, suppose age features can be robustly identified and separated from other facial features. In that case, it might be possible to develop techniques to remove age information from facial images while preserving utility for other facial analysis tasks. Understanding how age is encoded in facial features is a prerequisite for effectively managing age-related privacy. A robust age feature extractor, such as OrdCon, is vital for these privacy-preserving transformations.
V-A2 To other attributes
Based on Han et al. [90], facial attributes can be broadly categorized into ordinal and nominal attributes. Beyond age, OrdCon could be directly applicable to estimating other ordinal attributes, such as facial expression intensity [91] or the more precise facial action unit intensity [92]. While nominal attributes like gender and race lack ordinal scales, they are often entangled with ordinal attributes such as age or hair length [93, 94, 95]. In such cases, OrdCon could handle them similarly to how it addresses identity information in this work, i.e., by jointly learning both sets of features first and then reducing the variation in the nominal attribute caused by the ordinal attribute.
V-B Limitation
V-B1 Generalizability Evaluation Matrix
In this paper, we use two matrices, the alignment of positive samples and the divergence of direction vectors. The former is designed for categorical data and classification problems; hence, it does not apply to regression tasks with ordinal data. Also, we modified the divergence from measuring the averaged dot product of the class center to the direction vectors. The overall age progression direction in this work is defined as the direction from the youngest age to the eldest age, which may be inaccurate and not robust enough in cases where the distribution of age features does not follow a straight line. Therefore, more carefully designed evaluation matrices are needed for the analysis of the generalizability of contrastive learning frameworks on ordinal data.
V-B2 Homogeneous-dataset performance
OrdCon’s focus is primarily on extracting generalized features, which utilizes contrastive learning and order learning to model the aging progression in the feature space. That is the main reason it constantly outperforms other methods in cross-dataset experiments. However, its performance is surpassed by others’ in some homogeneous-dataset experiments. Specifically, MCGRL outperforms OrdCon in some homogeneous-dataset experiments on age estimation, and IEFP outperforms our framework in some homogeneous-dataset experiments on AIFR. Both methods focus on feature extraction that is effective at extracting dataset-specific features. Specifically, MCGRL uses graphs to learn facial structures, which is hard to generalize if the demographics change from dataset to dataset. On the other hand, IEFP leverages carefully designed purification units to learn the correlation between age and identity features. This relationship is also hard to generalize as aging progress is highly personalized. Although they are not designed for generalizability, OrdCon could learn from these methods by combining its contrastive learning objectives with customized feature modeling techniques to thrive in both cross-dataset and homogeneous-dataset settings.
VI Conclusion
@@
In this paper, we propose OrdCon to extract generalized age features for both age estimation and AIFR tasks. OrdCon is a new contrastive learning framework that leverages order learning to model natural age progression and a novel soft-proxy matching loss to minimize intra-class variance within each age cluster. The proposed method can be directly applied to age estimation. For AIFR, we employ a multitask feature extractor to simultaneously extract age and identity features, with a GRL used to make age information indistinguishable, thereby achieving age-invariant identity features. This paper also propose a new evaluation matrix to quantify the generalizability for regression methods. We demonstrate that our framework achieves either state-of-the-art or comparable performance in homogeneous-dataset experiments for both age estimation and the AIFR task. In addition, by utilizing contrastive learning objectives to improve generalizability, OrdCon outperforms other methods in cross-dataset experiments and generalizability evaluation matrices.
VII Acknowledgment